### v1

# SMITIN: Self-Monitored Inference-Time INtervention for Generative Music Transformers

Abstract
We introduce Self-Monitored Inference-Time INtervention (SMITIN), an approach for controlling an autoregressive generative music transformer using classifier probes. These simple logistic regression probes are trained on the output of each attention head in the transformer using a small dataset of audio examples both exhibiting and missing a specific musical trait (e.g., the presence/absence of drums, or real/synthetic music). We then steer the attention heads in the probe direction, ensuring the generative model output captures the desired musical trait. Additionally, we monitor the probe output to avoid adding an excessive amount of intervention into the autoregressive generation, which could lead to temporally incoherent music. We validate our results objectively and subjectively for both audio continuation and text-to-music applications, demonstrating the ability to add controls to large generative models for which retraining or even fine-tuning is impractical for most musicians.
Audio samples of the proposed intervention approach are available on our demo page.
Machine Learning, ICML
1 Introduction
Dynamic control of an audio signal can be considered a fundamental job of most audio professionals. Musicians must precisely control their instruments, and recording studio engineers manipulate various controls to achieve the desired result. Recent advances in the generative modeling of text and images have also been applied to audio signals, leading to an emerging body of literature on how to condition and control audio generative models (Dhariwal et al., 2020; Copet et al., 2023; Liu et al., 2023; Wu et al., 2023a; Garcia et al., 2023; Kreuk et al., 2022; Borsos et al., 2023).
Figure 1: Overall pipeline of SMITIN for inference-time intervention on a pre-trained music generative transformer. The process attempts to enforce specific musical factors (e.g., presence of a particular instrument) during the generation process. SMITIN utilizes a self-monitoring technique to dynamically adjust the intervention strength at each generation step, enabling precise control over the inclusion of the target characteristic while preserving the musical integrity of the output.
Among possible conditioning inputs, describing the audio a user wishes to generate in terms of natural language arguably provides the most flexibility. Such text-to-music models, however, lack the fine-grained temporal control often desired by audio professionals. This has led to training text-to-music models that allow for additional conditioning inputs, such as melody sequences (Copet et al., 2023), or other types of sequential input (Hawthorne et al., 2022; Donahue et al., 2023), but these models still suffer from the fact that, once trained, the type of conditioning input is fixed. To alleviate this limitation, some works have explored inference-time adaptation approaches that allow the addition of new types of control inputs without requiring retraining, i.e., allowing the users of text-to-music models to build their own controls, or “knobs” (Levy et al., 2023; Wu et al., 2023a; Novack et al., 2024; Lin et al., 2023, 2024). However, much previous work applies mainly to diffusion-based text-to-audio models, or it necessitates training supplementary adapters atop the foundational generative model to introduce additional control. To the best of our knowledge, no one has considered the application of inference-time control techniques to language model-like transformer-based text-to-audio models.
Audio “language models” typically work by first tokenizing chunks of an audio waveform, for example by passing them through the encoder of a pre-trained residual vector quantized autoencoder (Zeghidour et al., 2021; Défossez et al., 2022; Kumar et al., 2023), and then pass these audio tokens through an autoregressive transformer trained to predict the next token. Finally, a decoder converts the generated token into audio. Because this class of text-to-music models shares many architectural characteristics with large language models, it is interesting to explore whether techniques developed to provide inference-time control of LLMs (Subramani et al., 2022; Turner et al., 2023; Hernandez et al., 2023) may also be effective for audio. In particular, we take inspiration from recent work attempting to make text language models more truthful by taking advantage of the fact that the learned internal representations of these models are able to represent the concept of truthfulness (Li et al., 2023). Classifier probes (Alain & Bengio, 2016; Belinkov, 2022; Tenney et al., 2019) applied to each attention head are used to determine whether or not that attention head has learned to represent truthfulness, and inference-time intervention (ITI) is then used to surgically modify the outputs of only the most truth-correlated attention heads, resulting in an LLM that hallucinates less. Previous work (Castellon et al., 2021) has shown that classifier probes trained on the internal representations of generative music transformers can lead to strong performance across a wide variety of music classification tasks. However, these probes were trained using the learned representation from entire layers, not individual attention heads.
We first show using classifier probes that individual self-attention heads in a pre-trained state-of-the-art autoregressive music transformer, MusicGen (Copet et al., 2023), have indeed learned to represent aspects of music we may wish to build control knobs for, making it a priori suitable for ITI approaches such as in (Li et al., 2023). However, we find that direct application of this particular approach to a music transformer is sub-optimal because of the difficulty in generating long temporally coherent music samples, as too much intervention causes the generated music to quickly become incoherent. Furthermore, the method asks for the empirical tuning of the number of heads selected for intervention, impeding the ability to implement multiple intervention kinds at scale. To remedy this, we propose the inclusion of a self-monitoring process into the intervention operation, such that we only apply the intervention when the learned classifier probes tell us it is necessary based on the state of the generation network. Crucially, this self-monitoring technique enables real-time assessment of whether the current generated sample incorporates the target factor, allowing for the generation of musically aligned samples without a costly retraining or fine-tuning process. We also propose to weigh all heads as a function of their learned probe performance, removing the need for empirical head selection tuning.
2 Self-Monitored Inference-Time Intervention
2.1 Autoregressive Transformer Models
Architectures such as that of MusicGen currently form the foundation of state-of-the-art music generative models. These models are characterized by the autoregressive generation of a sequence of audio frames by transformer models. They include a collection of $L$ multi-head self-attention layers (residual connections, normalization layers, and fully-connected feed-forward layers are employed as usual and not described here). At current time step $t$ in the generated sequence, the $l$-th self-attention layer computes $H$ self-attention heads $z_{l,h}(t)\in\mathbb{R}^{D}$ from an input vector $x_{l}(t)\in\mathbb{R}^{DH}$ as
$z_{l,h}(t)=\text{Att}(W^{Q}_{l,h}x_{l}(t),W^{K}_{l,h}x_{l}(1\!:\!t),W^{V}_{l,h}x_{l}(1\!:\!t)),$ (1)
where $x_{l}(1\!:\!t)=[x_{l}(1),\dots,x_{l}(t)]$, $W^{Q}_{l,h}$, $W^{K}_{l,h}$, and $W^{Q}_{l,h}$ denote the head-specific query, key, and value projection matrices, all in $\mathbb{R}^{D\times DH}$, and Att denotes the attention operator (Vaswani et al., 2017). The output vector $y_{l}(t)\in\mathbb{R}^{DH}$ is obtained by projecting back each head into $DH$-dimensional space and summing:
$y_{l}(t)=\sum\limits^{H}_{h=1}W^{O}_{l,h}z_{l,h}(t),$ (2)
where $W^{O}_{l,h}\in\mathbb{R}^{DH\times D}$ is a projection matrix.
2.2 Inference-time Intervention (ITI)
(Li et al., 2023) first recognized that the output of the model could be somewhat controlled at inference by intervening in (i.e., modifying) the computation of the output of the multi-head self-attention layer in Eq. (2). This intervention is done by adding a term to the heads $z_{l,h}(t)\in\mathbb{R}^{D}$ before the projection:
$y_{l}(t)=\sum\limits^{H}_{h=1}W^{O}_{l,h}\left(z_{l,h}(t)+\alpha w_{l,h}\cdot\sigma_{l,h}\theta_{l,h}\right),$ (3)
where $\theta_{l,h}\in\mathbb{R}^{D}$ is a vector representing the head-specific ITI “direction” in the $D$-dimensional head-specific space, $\sigma_{l,h}\in\mathbb{R}^{+}$ is a head-specific normalization of $\theta_{l,h}$, $\alpha\in\mathbb{R}^{+}$ is the system-wide ITI strength, and $w_{l,h}\geq 0$ is a head-specific strength weighting. While the formulations are equivalent, we depart somewhat from (Li et al., 2023) by introducing $w_{l,h}$ as a separate term, whereas it was factored into $\theta_{l,h}$ in their notations. This helps to disambiguate the dual function played by $\theta_{l,h}$ in their formulation.
The head-specific ITI directions $\theta_{l,h}$ are obtained through classifier probes. Following (Li et al., 2023), we find a side probing dataset suited to train a classifier mirroring our ITI goal (e.g., drum classifier if the goal is to add drums). We then run the dataset through our (frozen) generation network and collect a set of heads $z_{l,h}(t)$ for all $l,h$. For each $l$ and $h$, we then train a distinct logistic regression classifier probe with parameters $\tilde{\theta}_{l,h}$, whose prediction is obtained as $\text{sigmoid}(\langle\tilde{\theta}_{l,h},z_{l,h}(t)\rangle)$. Once training is complete, we set ITI direction $\theta_{l,h}$ as the final $\tilde{\theta}_{l,h}$, and $\sigma_{l,h}$ as the standard deviation of $\langle\theta_{l,h},z_{l,h}(t)\rangle$ for all $z_{l,h}(t)$ obtained on the combined probing training and testing data. We also take note of the final classifier accuracy $\text{acc}_{l,h}$ on the probing test data. In the case of MusicGen, we apply the generative model in unconditional generation mode for probing, as our probing dataset generally lacks text queries for its audio samples.
In (Li et al., 2023), head-specific weights $w_{l,h}$ are set by finding the top-$K$ heads in terms of classifier probe accuracy $\text{acc}_{l,h}$, setting their $w_{l,h}$ to 1 and the others to 0. The most effective $\alpha$ and $K$ are found by hyperparameter grid search.
2.3 Sparse intervention
In music generation, the approach above may present some limitations. For example, for the task of audio continuation, we observe that ITI often leads to changes that are too abrupt to be musically plausible (see Section 3.3). As mitigation, we propose to diminish the ITI frequency across time steps, potentially allowing the generation process to better align with the underlying rhythmic structure of the generated music. This corresponds to replacing the ITI weights $w_{l,h}$ in Eq. (3) by time-varying weights $w_{l,h}(t)$, which can only be non-zero for $t=t_{0}+is,$ $i\in\mathbb{N}$, where $t_{0}$ is an intervention start time and $s$ represents the number of steps between each ITI (e.g., $s=5$ to perform ITI every 5 time steps). The value of $w_{l,h}(t)$ for $t=t_{0}+is$ can be set by another criterion, such as the original one based on top-$K$ heads.
2.4 Soft-weighting
One crucial limitation of (Li et al., 2023) is to leave the number $K$ of heads on which ITI is performed (i.e., the number of pairs $(l,h)$ for which $w_{l,h}(t)\neq 0$) as hyperparameter to tune. We additionally propose a hyperparameter-free soft-weighting approach based on the collected probe accuracies $\text{acc}_{l,h}$, and show that it is sufficient to perform effective ITI (see Tab. 4.1). In practice, we propose setting the weights as
$w_{l,h}(t)=\left(\frac{\text{acc}_{l,h}-\text{acc}_{\text{min}}}{\text{acc}_{\text{max}}-\text{acc}_{\text{min}}}\right)^{\!c},$ (4)
with $\text{acc}_{\text{min}}$ and $\text{acc}_{\text{max}}$ the minimum and maximum accuracies recorded across all $l$ and $h$, and $c$ a power factor allowing to modulate the relative weights of heads with accurate vs. inaccurate classifier probes (we use $c=3$). By construction, $w_{l,h}(t)$ is guaranteed to fall between 0 and 1.
2.5 Automated Intervention Modulation by Self-monitoring
Ideally, we would expect a system capable of modulating the ITI strength to be most effective, as systems with time-invariant $w_{l,h}(t)$ make no use of (and, as such, cannot react to) the state of the inference model during generation.
As our core contribution, we propose to use the classifier probes to drive such a modulation. We first define as $\mathcal{H}_{K}$ the set of top-$K$ heads $(l,h)$ by probe classifier accuracy. For each new generation time step, we collect the set $\mathcal{C}(t)$ of output predictions of the classifier probes for all heads in $\mathcal{H}_{K}$ before intervening on them, i.e.,
$\mathcal{C}(t)=\left\{\text{sigmoid}\left(\langle\theta_{l,h},z_{l,h}(t)\rangle\right)\,|\,(l,h)\in\mathcal{H}_{K}\right\}.$ (5)
The statistics of $\mathcal{C}(t)$ reflect the confidence of the probes regarding the success of the intervention. By comparing them to the baseline accuracies obtained on the probing training data, we can devise a modulation scheme to update $w_{l,h}(t)$. We compute the median $\bar{\mathcal{C}}(t)=\text{med}(\mathcal{C}(t))$ and the change in this median since the prior generation step with intervention, $\Delta(t)=\bar{\mathcal{C}}(t)-\bar{\mathcal{C}}(t\!-\!s)$. We then define a threshold based on the median and standard deviation of the set $\mathcal{A}$ of probe accuracies $\text{acc}_{l,h}$ on the probing training data over the same heads $\mathcal{H}_{K}$,
$\mathcal{A}=\{\text{acc}_{l,h}\,|\,(l,h)\in\mathcal{H}_{K}\},$ (6)
setting the threshold as $\tau=\text{med}(\mathcal{A})-\text{std}(\mathcal{A})$.
The update algorithm is then defined following:
  • For the first generation time step with ITI $t_{0}$, $w_{l,h}(t_{0})$ is set following Eq. (4), and we set $\Delta(t_{0})\leftarrow 0$,
  • For each subsequent generation time step with ITI $t\!+\!s$, we have 3 cases:
    1. 1.
      if $\bar{\mathcal{C}}(t)\!<\!\tau$ and $w_{l,h}(t)\!>\!0$, we set
      $w_{l,h}(t+s)\leftarrow w_{l,h}(t)\cdot(1-\Delta(t)),$ (7)
    2. 2.
      if $\bar{\mathcal{C}}(t)\!<\!\tau$ but $w_{l,h}(t)=0$, we reset to the initial value
      $w_{l,h}(t+s)\leftarrow w_{l,h}(t_{0}),$ (8)
    3. 3.
      if $\bar{\mathcal{C}}(t)\geq\tau$, we set $w_{l,h}(t+s)\leftarrow 0$.
3 Experiments
Here, we validate the effectiveness of our proposed contribution by presenting a thorough investigation of ITI applied to the music generative task of instrument addition. This section outlines methods to quantify the effectiveness of ITI techniques objectively, complemented by a subsequent subjective evaluation.
3.1 Setup
For the evaluation of ITI on music transformer, we utilize the large mono configuration, which is composed of 48 layers with 32 attention heads each, amounting to a total of 1,536 heads. Unless otherwise noted, the self-monitoring configuration of SMITIN is set to an intervention strength of $\alpha=5.0$, a sparse intervention rate of $s=5$, and $K=16$ for the selection of monitoring probes.
To enable the probing mirroring the instrument addition task, we curate probing datasets for drums, bass, guitar, and piano detection from MusDB (Rafii et al., 2019) and MoisesDB (Pereira et al., 2023), both of which provide multi-track recordings with isolated instrument stems. Stems from their training (resp. testing) partitions are used to form the probing training (resp. testing) partitions. To create positive and negative classes, i.e., one including and one excluding the target instrument, we proceed as follows. We collect a first set consisting of the mixed audio from each track recording after removing the target instrument stem, and trim silent sections. We then collect a second set consisting of the target instrument stems, and trim silent sections. We then generate negative samples using audio segments randomly sampled from the first set, and positive samples by randomly mixing a mixture without the target instrument (from the first set) with a stem of the target instrument (from the second set). Following silence trimming, the datasets amount to 8.3, 8.0, 13.1, and 5.7 hours for drums, bass, guitar, and piano, respectively. These durations represent the total amount of paired data available for probing. Since the negative samples are more abundant, we balanced the dataset to ensure an equal number of positive and negative samples, based on the total duration of positive samples.
3.1.1 Generation Approach
Our experimental setup includes two contexts in which to perform ITI: audio continuation and Text-to-Music. In both cases, the objective is to perform the task while adding the target instruments (drums, bass, guitar, and piano) into the music pieces through ITI.
Audio Continuation: Given an input music signal, we continue to generate a music piece while trying to add the target instrument. The input segment is a 3-second-long music segment obtained from the MusDB and MoisesDB test datasets, which does not contain the target instrument.
Text-to-Music: This task uses text prompts from the MusicCaps (Agostinelli et al., 2023) dataset to guide the generation process, allowing us to evaluate the effectiveness of ITI in the context of text-conditioned generation. We use the aspect list (e.g., “pop, tinny wide hi hats, mellow piano melody”) over the free text caption to further investigate the impact of adding simple instruction “, add inst.” to the list.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 5: Inferred accuracy of the top-K probes’ monitored decision along the time-axis. The yellow line, green line, and the shaded blue region denote the median, mean, and standard deviation of inferred outputs by the probes, respectively. (Left) Monitored result on a real-world music sample. The high accuracy (close to 1.0) until 3.5 seconds reflects the actual presence of drums, which aligns with the audio sample where drums are present only up to that point. (Right) The four sub-plots display the results of audio continuation on the same input music with varying ITI frequencies ($s=[1,5,10,20]$). These illustrate that more frequent intervention leads to a swifter convergence towards the target musical factor, at the expense of losing musical consistency with the input music.
3.1.2 Objective Metrics
Our objective evaluation comprises three metrics designed to assess the effectiveness of the ITI process. We generate 1,000 music pieces, each 30 seconds long, for each experimental configuration.
Success Rate: We introduce Success Rate to quantify the efficacy of ITI in successfully adding the desired instrument to the music. This metric is derived from our self-monitoring technique from Section 2.5, which measures the likelihood of the target musical factor based on the timewise probability inferred by the top-$K$ best-performing probes, where $K=16$. For each audio, the success rate is calculated as $n_{\text{success}}/N$, where $n_{\text{success}}$ is the number of samples $t$ after the start of the intervention for which $\bar{\mathcal{C}}(t)>\tau$, and $N$ is the total number of time steps generated after the intervention. The reliability of this metric is contingent upon the top-$K$ probes’ test performance in the probing task.
Fréchet Audio Distance (FAD): FAD (Kilgour et al., 2018) is used to compare generated music with real music datasets, which helps ensure that the intervention does not cause an unrealistic shift in MusicGen’s output distribution. Instead of the conventional approach of using VGGish (Hershey et al., 2017) as the deep encoder, we use L-CLAP mus (Wu et al., 2023b) for dimension reduction, as it has been shown to offer a more accurate representation of music (Gui et al., 2023). As the real music datasets, we adopt MusDB or MoisesDB for audio continuation and MusicCaps for Text-to-Music generation.
Similarity Measurement in Audio Continuation: For the audio continuation task, we measure the musical similarity between input and generated samples. We employ the Music Effects Encoder (Koo et al., 2022), a model trained with self-supervised contrastive learning designed to encode song identity. The encoder is used to embed both input and generated samples, and their cosine similarity is calculated. This metric assesses how well the identity of the input music is preserved in the generated output.
3.2 MusicGen Probing
[…]
Table 1: Objective evaluation on Audio Continuation. FADL-CLAP ${}_{\text{L-CLAP mus}}$ is computed with MusDB (drums/bass) and MoisesDB (guitar/piano).
Method Success Rate [%] ($\uparrow$) FADL-CLAP ${}_{\text{L-CLAP mus}}$ ($\downarrow$) Similarity ($\uparrow$)
ITI Configuration drums bass guitar piano avg. drums bass guitar piano avg. drums bass guitar piano avg.
X unconditioned $12.5$ $13.5$ $3.6$ $0.0$ $7.4$ $0.326$ $0.253$ $0.377$ $0.320$ $0.319$ $0.864$ $0.933$ $0.929$ $0.941$ $0.916$
“add <inst.>” $16.6$ $27.4$ $5.5$ $1.7$ $12.8$ $0.364$ $0.301$ $0.385$ $0.351$ $0.350$ $0.833$ $0.905$ $0.929$ $0.936$ $0.900$
[…]
weight decay ITI $13.2$ $17.5$ $15.4$ $3.1$ $12.3$ $0.336$ $0.259$ $0.378$ $0.337$ $0.327$ $0.871$ $0.927$ $0.933$ $0.941$ $0.918$
SMITIN $20.6$ $30.4$ $33.8$ $8.7$ $23.3$ $0.346$ $0.267$ $0.397$ $0.337$ $0.336$ $0.857$ $0.922$ $0.935$ $0.938$ $0.913$
SMITIN + “add <inst.>” $16.9$ $28.0$ $16.9$ $17.4$ $19.8$ $0.381$ $0.303$ $0.391$ $0.374$ $0.362$ $0.839$ $0.909$ $0.929$ $0.942$ $0.904$
Table 2: Objective evaluation on Text-to-Music. FADL-CLAP ${}_{\text{L-CLAP mus}}$ is computed with MusicCaps.
Method Success Rate [%] ($\uparrow$) FADL-CLAP ${}_{\text{L-CLAP mus}}$ ($\downarrow$)
ITI Configuration drums bass guitar piano avg. drums bass guitar piano avg.
X text $29.9$ $50.4$ $17.4$ $7.1$ $26.2$ 0.482 (indep. of inst.)
text + “add <inst.>” $34.4$ $56.1$ $22.5$ $10.5$ $30.8$ $0.474$ $0.481$ $0.473$ $0.488$ $0.479$
[…]
text + SMITIN ($\alpha=5.0$) $29.9$ $53.4$ $23.1$ $12.5$ $29.7$ $0.471$ $0.493$ $0.499$ $0.487$ $0.487$
text + SMITIN ($\alpha=10.0$) $40.9$ $60.7$ $35.4$ $21.0$ $39.5$ $0.481$ $0.516$ $0.507$ $0.485$ $0.497$
3.3 Qualitative Evaluation
We qualitatively analyze the performance of top-$K$ probes’ monitoring capability through Fig. 5. The red dashed line in each plot indicates the threshold value $\tau$, as mentioned in Section 2.5. From the left side of the figure, it is apparent that the median (yellow line) prediction more closely aligns with the actual presence of the target instrument than the mean prediction (green line). Consequently, we use $\bar{\mathcal{C}}(t)=\text{med}(\mathcal{C}(t))$ to monitor during inference time.
The sub-plots on the right illustrate how the hyper-parameter $s$ influences the outcome of the audio continuation task. We observe a relationship between the intervention frequency $s$ and the rapidity with which the model aligns toward the target factor. This suggests that more frequent interventions cause the transformer to prioritize the target factor more heavily in subsequent token generation, potentially at the expense of referencing previous musical context. We discuss this trade-off between intervention quantity and musical coherence throughout this section.
3.4 Objective Evaluation
Utilizing the metrics introduced in Section 3.1.2, we objectively evaluate our methods on Audio Continuation and Text-to-Music. We compare with baseline methods that do not incorporate ITI, such as MusicGen’s standard unconditional and text-prompt conditioned generations. Additionally, we benchmark against the original ITI method, which applies a uniform intervention at every time step (i.e., $s=1$) with a configuration of $\alpha=5.0$ and $K=64$ to apply the intervention with the equal amount $\alpha$ to each of the $K$ selected heads. For Audio Continuation, we examine another ITI approach, weight decay, where $\alpha$ is progressively reduced across time steps during the generation process. For Text-to-Music, our evaluation includes both our default configuration ($\alpha=5.0$) and a stronger intervention level of $\alpha=10.0$ to observe the trade-off between the intervention strength and the preservation of musical fluency within the generated samples.
Table 3.2 presents the objective evaluation for the Audio Continuation task. The baseline methods reveal that MusicGen inherently tends to add certain instruments without explicit direction with varying success rates across instruments; drums and bass are more easily added compared to guitar and piano. Intriguingly, text-prompted generation does not consistently achieve the target instruction. Intervention methods, on the other hand, consistently outperform baseline approaches in successfully directing MusicGen toward adding the desired instruments, as indicated by the success rate. The original ITI, applying a constant level of intervention at each generation step, leads to a higher success rate but significantly alters the distribution of generated content, which can be observed via the similarity score for maintaining musical consistency. In contrast, the weight decay ITI approach has a subtler influence on the quality but produces a lower success rate. Our proposed SMITIN shows a notable 10.5% jump over text-prompt conditioning, and is better at retaining the model’s output distribution and generating consistent music. When SMITIN is deployed in conjunction with text-conditioning, it outperforms text-based instructions, but not to the same extent as when used independently. This result opens up the potential for integrating ITI with textual directives, allowing a dual axis of control—semantic guidance via text and fine-tuned adjustments through ITI.
For the Text-to-Music task, we explore the interplay of ITI with text prompts, as shown in Table 3.2. In this context, we utilize MusicCaps text prompts as the basis of MusicGen’s performance. Consistent with earlier observations, the original ITI method, despite showing the highest success rate in instrument addition, performs the lowest in terms of FAD, indicating a more pronounced shift from the natural music distribution. SMITIN, in its base configuration ($\alpha=5.0$), achieves an average performance on par with additional text guidance. We further analyze SMITIN with an increased intervention strength of $\alpha=10.0$, and observe that a clear trade-off emerges between the success rate and the naturalness of the generated music. This outcome highlights the flexibility of SMITIN, offering users a tunable “knob” to balance between precision in achieving specific musical characteristics and maintaining the authenticity of the musical piece.
3.5 Subjective Evaluation
Table 3: Average and 95% confidence interval of subjective listening tests for overall audio quality across four instrument addition interventions for audio continuation and text-to-music experiments.
Method Continuation Text-to-Music
No intervention $3.35$$\pm$ $0.10$ $3.67$$\pm$ $0.09$
“add <inst.>” $3.27$$\pm$ $0.10$ $3.61$$\pm$ $0.09$
original ITI $3.28$$\pm$ $0.11$ $3.55$$\pm$ $0.09$
SMITIN $3.34$$\pm$ $0.10$ $3.80$$\pm$ $0.09$
MoisesDB 3.98 $\pm$ 0.29
To further validate the effectiveness of our intervention techniques on MusicGen, we performed a subjective listening test through Amazon Mechanical Turk following the best practices from (Ribeiro et al., 2011). The purpose of this test was not to evaluate the quality and relevance of MusicGen as this was already explored in (Copet et al., 2023), but rather to provide quantitative evidence that: (1) our proposed success rate objective metric correlates with human perception, and (2) our proposed intervention techniques are not detrimental to overall audio quality. To do this, we assessed whether human listeners could detect the presence of the instruments added by our intervention technique, and asked them to rate samples in terms of overall quality on a scale of 1-5. We also included real music samples from MoisesDB to provide a performance baseline. Table 3 compares the overall audio quality ratings, where we note that our proposed intervention techniques perform comparably to applying no intervention. Furthermore, we computed the Spearman correlation coefficient between the objective success rate metric and the average score across human raters asked to detect the presence of the instrument added by the intervention techniques. Across all intervention algorithms, we achieved $\rho=0.48$ for audio continuation and $\rho=0.67$ for text-to-music, both with p-values much less than 0.05, indicating that our proposed success rate metric correlates with the ability of humans to detect instruments added through intervention. Further results and details on the experimental setup are provided in Appendix A.
Table 4: Ablation study on multi-directional ITI (performance averaged over all instruments)
Success Rate [%]
Method indiv. simult. FADL-CLAP ${}_{\text{L-CLAP mus}}$ Similarity
unconditioned $11.1$ $0.1$ $0.419$ $0.781$
text $27.8$ $0.43$ $0.369$ $0.664$
SMITIN $25.3$ $1.23$ $0.445$ $0.734$
[…]
4 Analysis
4.1 Intervention with Multiple Directions
We explore the application of ITI in scenarios where multiple musical elements are introduced simultaneously. In our setup, we focus on the continuation of music that has one instrument present and assess the addition of three others. For instance, the intervention task is to generate drums, bass, and piano when the input music has only guitar. We consider two types of success rates, an individual one for each instrument as before, and a simultaneous one to assess how often all 3 target instruments were added together. We consider a successful case for a song if each instrument is deemed present more than 50 % of the time.
The summarized results in Table 4 reveal that while text prompts yield higher success rates for individual instruments, their generated output diverges from the input music’s characteristics. This indicates that text conditioning may not adequately consider the input and opts to generate new content based solely on the given instruction. Without ITI, MusicGen tends to produce continuations focused on a single instrument, which deviates from the desired complex mixtures, as evidenced by the FAD scores. SMITIN, however, significantly outperforms text prompts in preserving input music similarity, and further enables more granular control over each musical aspect, leading to a higher success rate of generating all desired instruments in unison. This fine-grained control mechanism bolsters SMITIN’s potential as a robust tool for complex music generation tasks where maintaining the essence of the input is crucial.
Table 5: Ablation study: top-$K$ head selection vs. head soft-weighting
Top-K Interventions Success Rate [%] ($\uparrow$) FADL-CLAP ${}_{\text{L-CLAP mus}}$ ($\downarrow$) Similarity ($\uparrow$)
drums bass guitar piano avg. drums bass guitar piano avg. drums bass guitar piano avg.
K = 1600 $14.9$ $24.8$ $11.9$ $3.9$ $13.8$ $0.327$ 0.249 0.392 $0.335$ 0.325 $0.864$ 0.929 $0.931$ $0.941$ $0.916$
K = 3200 $14.4$ $20.2$ $16.9$ $5.6$ $14.2$ 0.325 $0.259$ $0.393$ $0.334$ $0.327$ $0.867$ $0.928$ $0.934$ $0.941$ 0.917
K = 6400 $15.3$ $26.5$ $17.4$ $11.0$ $17.5$ $0.335$ $0.267$ $0.390$ $0.336$ $0.332$ 0.869 $0.923$ $0.934$ $0.942$ 0.917
K = 1280 $18.2$ $33.6$ $27.4$ $13.8$ $23.2$ $0.352$ $0.277$ $0.411$ 0.327 $0.341$ $0.860$ $0.924$ $0.932$ 0.944 $0.915$
K = 2560 $23.9$ $36.0$ $36.8$ $16.8$ $28.3$ $0.370$ $0.283$ $0.405$ $0.332$ $0.347$ $0.848$ $0.920$ $0.933$ 0.944 $0.911$
K = 5120 $30.0$ $50.5$ $47.8$ $29.7$ $39.5$ $0.386$ $0.300$ $0.411$ $0.339$ $0.359$ $0.837$ $0.917$ $0.932$ $0.942$ $0.907$
K = 1024 32.9 79.0 61.1 43.5 54.1 $0.560$ $0.355$ $0.417$ $0.363$ $0.423$ $0.823$ $0.895$ $0.930$ $0.936$ $0.896$
[…]
4.2 Generating “Realistic” Music
To generate more realistic music through ITI, we fit probes to discern between real music (DISCO-10M dataset’s DISCO-200K-high-quality subset (Lanzendörfer et al., 2023)) and synthetic outputs from MusicGen itself. Surprisingly, the probing showed high performance with the best-performing probe achieving a 96.2% accuracy rate and an average accuracy of 77.5% in distinguishing real from synthetic music (See probing results in Appendix LABEL:fig:probe_acc_real_fake). Leveraging this result, we then steer MusicGen in an audio continuation task, using real-world music inputs from the DISCO-10M dataset to generate continuations that maintain the “realistic” quality.
Analyzing the quantitative results depicted in Fig. 6, simple text prompts prove insufficient in preserving the realistic nature of the music; however, ITI emerges as a more effective approach. Regarding FAD scores, configurations of SMITIN with $\alpha$ values of 1 and 5 yield smaller deviations in distribution compared to text-based methods. Even with an increased intervention strength ($\alpha=10$), SMITIN does not shift the distribution more than that of the text prompt “high-quality realistic music.”
Table 6: Ablation study on the number $n$ of paired data for probing (performance is averaged over all instruments, shows full-set success rate).
Number of Paired Data Success Rate [%] FADL-CLAP ${}_{\text{L-CLAP mus}}$ Similarity Probes Max. Acc. [%]
$n$ = 10 $31.0$ $0.357$ $0.909$ 79.1
$n$ = 100 $31.8$ $0.349$ $0.914$ 82.6
$n$ = 500 $27.3$ $0.341$ $0.915$ 84.4
$n$ = 1000 $25.5$ $0.340$ $0.914$ 84.8
[…]
4.3 Effects on Number of Probing Data
To investigate the influence of probing data quantity on SMITIN, we test the system with varying sizes of probing datasets randomly selected from the complete set. These subsets consist of $n=10$, $100$, $500$, and $1000$ data pairs, equivalent to music durations of $1$, $10$, $50$, and $100$ minutes respectively. The results are presented in Table 4.2, where we show the “full-set success rate” which is computed using the classifier probes trained with the full dataset every time.
The results indicate that the system demonstrates reasonable results even with a minimal dataset of just 1 minute. As the data size increases to $n=500$, the objective metrics stabilize, indicating an optimal data quantity threshold for effective probe training. An interesting observation from our results is the higher full-set success rates associated with smaller datasets. This trend suggests that probes trained on limited data may have lower confidence in their decisions, prompting them to recommend more interventions during the ITI process to meet the success threshold. At the same time, we observe a clear correlation between better probes (represented here through the accuracy of the best probe) and better ITI performance, whether measured as higher full-set success rate or similarity, or lower FAD.
This altogether has practical implications for the application of ITI in real-world scenarios. It further supports the extensive use of probing as the basis for effective ITI design. But it also suggests that users can effectively leverage ITI with only a small amount of their data to control generation. It means that effective ITI can be accessible and achievable without the need for extensive datasets, which is particularly beneficial for individual artists or smaller studios. This accessibility broadens the potential for creative and personalized applications of ITI in music generation, making it a versatile tool for a wide range of users.
4.4 Soft-weighting
This ablation study justifies the soft-weighting technique over selecting top-$K$ heads for ITI. The rationale for head-weighting arises from the observation that the optimal set of top-$K$ heads varies depending on the task or instrument. By adopting soft-weighting, we can bypass the need for hyper-parameter tuning specific to each task, simplifying the ITI process.
To validate this approach, we compare the performance of the full SMITIN (with soft-weigthing) against SMITIN where the initial weights $w_{l,h}(t_{0})$ are instead 1 for the top-$K$ most accurate heads according to probing (i.e., $(l,h)\in\mathcal{H}_{K}$), and 0 otherwise. Results from Table 4.1 indicate that while top-$K$ selection can yield good performance, it often requires task-specific tuning of $K$. In contrast, soft-weighting demonstrates a well-balanced performance across various metrics without necessitating such tuning. This balance is crucial because it means soft-weighting can adapt to various musical tasks and preferences, making ITI more flexible and user-friendly. However, it is important to acknowledge that soft-weighting may not always be the optimal choice for every ITI application. Users might prefer the top-$K$ approach depending on their specific needs and the nature of their musical goals.
5 Conclusions and Future Work
We proposed a novel approach for inference-time control of generative music transformers, which self-monitors probe accuracy to impose desired musical traits while maintaining overall music quality. A limitation of our approach is the dependence of performance on probe accuracy, that is if the underlying pre-trained transformer has not accurately learned a concept, or if the probe training set is inadequate, the success rate of our intervention technique may suffer (although it still could be creatively useful). In the future, we plan to explore additional intervention using additional musical traits such as genre, emotion, etc. Furthermore, we hope to investigate building “knobs” for interactively controlling the hyper-parameters necessary for inference-time control, in an effort to enable new technical and creative interactions.
Appendix A Subjective Listening Test Setup and Detailed Results
In selecting files for the listening test discussed in Section 3.5, for each algorithm and instrument addition intervention, we randomly select 20 files above the median objective success rate, and 20 files below the median. This leads to 1280 generated files to be rated (20 files x 2 top/bottom x 4 instruments x 4 algorithms x 2 audio continuation/text-to music). We additionally include real music samples from MoisesDB as references. All audio files included in the listening test were ten seconds in length and normalized to -12 loudness units full scale (LUFS) (Grimm et al., 2010). We obtain at least 3 ratings per file. For the audio continuation experiments, we ask raters to ignore the first 3 seconds of each audio file, as that conditioning signal intentionally does not include the target instrument.
Table 7 displays the overall objective quality results broken out by instrument and algorithm. In general, the performance of all algorithms are quite similar, with SMITIN an no intervention having the best overall quality.
In addition to overall quality, we also ask whether listeners can detect the presence of the instruments targeted by our intervention. This serves to help validate our success rate objective metric. Table 8 compares the Spearman correlation coefficient $\rho$ computed between the success rate objective metric and average human rating computed across all processing algorithms. All correlation coefficients in Table 8 have p-value less than 4e-5. We also compare the average human score with the average success rate objective metric (Alg.) for the top-ranked and bottom-ranked files selected for the listening test. For the top ranked files the scores appear to match quite well, while for the bottom ranked files, it seems human raters tend to overestimate the presence of most instruments compared to the objective metric. We also hypothesize that the lower match for audio continuation in Table 8 compared to text to music may be due in part to the test being more subjectively difficult as raters have to focus on the end of the audio file under test.
Table 7: Average and 95% confidence interval of subjective listening tests for overall audio quality across four instrument addition interventions for both audio continuation and text-to-music experiments.
Method Audio Continuation Text-to-music
drums bass guitar piano avg drums bass guitar piano avg
No intervention $3.22$$\pm$ $0.21$ $3.75$$\pm$ $0.18$ $3.26$$\pm$ $0.22$ $3.16$$\pm$ $0.18$ $3.35$$\pm$ $0.10$ $3.63$$\pm$ $0.18$ $3.66$$\pm$ $0.20$ $3.70$$\pm$ $0.19$ $3.70$$\pm$ $0.17$ $3.67$$\pm$ $0.09$
“add <inst.>” $3.31$$\pm$ $0.21$ $3.41$$\pm$ $0.18$ $3.06$$\pm$ $0.211$ $3.31$$\pm$ $0.22$ $3.27$$\pm$ $0.10$ $3.78$$\pm$ $0.18$ $3.61$$\pm$ $0.18$ $3.54$$\pm$ $0.21$ $3.50$$\pm$ $0.19$ $3.61$$\pm$ $0.09$
original ITI $3.42$$\pm$ $0.22$ $3.04$$\pm$ $0.21$ $3.54$$\pm$ $0.20$ $3.12$$\pm$ $0.22$ $3.28$$\pm$ $0.11$ $3.62$$\pm$ $0.19$ $3.61$$\pm$ $0.18$ $3.65$$\pm$ $0.19$ $3.33$$\pm$ $0.20$ $3.55$$\pm$ $0.09$
SMITIN $3.40$$\pm$ $0.20$ $3.45$$\pm$ $0.18$ $3.27$$\pm$ $0.21$ $3.23$$\pm$ $0.19$ $3.34$$\pm$ $0.10$ $3.93$$\pm$ $0.18$ $3.75$$\pm$ $0.17$ $3.77$$\pm$ $0.18$ $3.73$$\pm$ $0.17$ $3.80$$\pm$ $0.09$
MoisesDB 3.98 $\pm$ 0.29
Table 8: Subjective results ranking presence of instruments added by intervention, across multiple algorithms.
Audio Continuation Text-to-music
Top Ranked Bottom Ranked Top Ranked Bottom Ranked
Instrument $\rho$ Human Alg. Human Alg. $\rho$ Human Alg. Human Alg.
drums $0.76$ $0.88$ $0.81$ $0.19$ $0.13$ $0.80$ $0.95$ $0.95$ $0.16$ $0.18$
bass $0.42$ $0.78$ $0.89$ $0.48$ $0.06$ $0.58$ $0.80$ $0.96$ $0.38$ $0.13$
guitar $0.40$ $0.78$ $0.76$ $0.52$ $0.15$ $0.67$ $0.86$ $0.88$ $0.34$ $0.14$
piano $0.32$ $0.60$ $0.72$ $0.33$ $0.32$ $0.55$ $0.70$ $0.79$ $0.31$ $0.23$
avg. $0.48$ $0.76$ $0.80$ $0.39$ $0.17$ $0.67$ $0.83$ $0.90$ $0.30$ $0.17$
Appendix B Objective Evaluation
B.1 Instrument Removal
We explore an additional application: removing the target instrument during Audio Continuation using the same probes employed for instrument addition. To accomplish this, we apply negative intervention strength $\alpha$ to eliminate the target instrument from given input music. The hyperparameter settings for this task are configured as $\alpha=-10.0$ for both original ITI and SMITIN, with $s=1$ and $\tau=0.5$ for SMITIN. The success rate in this context is defined as $\bar{\mathcal{C}}(t)<\tau$, where $\tau=0.5$.
Table B.1 reveals an inverse tendency for each instrument’s removal: drums and bass tend to be continuously generated, whereas guitar and piano are more likely to be omitted during the generation process. Moreover, we find that both text-conditioning prompts: “remove <inst.>”, and “no <inst.>” are ineffective in removing the target instrument. This outcome suggests another aspect of controllability that is challenging to achieve with text prompts alone.
Table 9: Objective evaluation on Instrument Removal.
Method Success Rate [%] ($\uparrow$) FADL-CLAP ${}_{\text{L-CLAP mus}}$ ($\downarrow$) Similarity ($\uparrow$)
ITI Configuration drums bass guitar piano avg. drums bass guitar piano avg. drums bass guitar piano avg.
X unconditioned $2.2$ $16.2$ $14.1$ $31.6$ $16.0$ $0.442$ $0.517$ $0.338$ $0.513$ $0.452$ $0.884$ $0.579$ $0.847$ $0.816$ $0.781$
“remove <inst.>” $1.8$ $4.8$ $19.2$ $23.2$ $12.2$ $0.568$ $0.634$ $0.420$ $0.789$ $0.602$ $0.871$ $0.545$ $0.795$ $0.808$ $0.754$
“no <inst.>” $7.3$ $10.8$ $17.0$ $32.5$ $16.9$ $0.267$ $0.267$ $0.307$ $0.289$ $0.282$ $0.934$ $0.889$ $0.937$ $0.943$ $0.925$
[…]
SMITIN $9.1$ $5.1$ $79.9$ $78.4$ $43.1$ $0.481$ $0.538$ $0.481$ $0.455$ $0.488$ $0.779$ $0.545$ $0.691$ $0.778$ $0.698$
B.2 Intervention with Different Directions
Following (Li et al., 2023), we compare intervention results using directions as logistic regression classifiers’ weights $\tilde{\theta}_{l,h}$ and mass mean shift. The intervention direction for mass mean shift is determined during probing as a vector that directs from the centroid of negative activations to the centroid of positive activations. The results are shown in Table 10. Similar to (Li et al., 2023), we observe that using mass mean shift as the intervention direction better achieves the target musical factor (according to success rate). However, we again observe a trade-off between controllability and generation distribution shift. Considering that generation quality is crucial in the music generation task, we adopt $\tilde{\theta}_{l,h}$ as the final intervention direction.
Table 10: Comparison with different intervention directions on Audio Continuation.
Success Rate [%] ($\uparrow$) FADL-CLAP ${}_{\text{L-CLAP mus}}$ ($\downarrow$) Similarity ($\uparrow$)
ITI Direction drums bass guitar piano avg. drums bass guitar piano avg. drums bass guitar piano avg.
$\tilde{\theta}_{l,h}$ $20.6$ $30.4$ $33.8$ $8.7$ $23.3$ 0.346 $0.267$ $0.397$ $0.337$ 0.336 $0.857$ $0.922$ $0.935$ $0.938$ 0.913
mass mean shift $17.0$ $48.6$ $36.0$ $9.7$ 27.8 0.355 $0.291$ $0.398$ $0.323$ $0.341$ $0.855$ $0.923$ $0.928$ $0.941$ $0.911$
B.3 FAD Compared with Unconditioned Generation
To support the claim that SMITIN does not significantly alter the original distribution of MusicGen, we compute the FAD score using the distribution of unconditioned generation music outputs. For Text-to-Music, the comparison distribution is that of text-conditioned generation using MusicCaps text prompts, as discussed in Table 3.2. According to Table B.3, the FAD score of “add <inst.>” is significantly lower in Text-to-Music than in Audio Continuation, indicating that the output distribution between unconditional and conditional generation experiences a significant shift. We observe that SMITIN with $\alpha=5.0$ exhibits the closest distribution distance to the original MusicGen. This supports the claim that SMITIN achieves the desired controllability while retaining MusicGen’s generation capability.
Table 11: FAD compared with the distribution of unconditioned generation on Audio Continuation and Text-to-Music.
FADL-CLAP ${}_{\text{L-CLAP mus}}$($\downarrow$) w/. unconditioned
Method Audio Continuation Text-to-Music
ITI Configuration drums bass guitar piano avg. drums bass guitar piano avg.
X “add <inst.>” $0.042$ $0.026$ $0.026$ $0.064$ $0.03$ $0.013$ $0.009$ $0.011$ $0.009$ $0.01$
[…]
SMITIN ($\alpha=5.0$) $0.022$ $0.017$ $0.018$ $0.010$ 0.016 $0.012$ $0.009$ $0.010$ $0.006$ 0.009
SMITIN ($\alpha=10.0$) $0.079$ $0.042$ $0.033$ $0.014$ $0.042$ $0.045$ $0.017$ $0.019$ $0.008$ $0.022$
B.4 Detailed Results of Ablation Studies
We disclose objective results on ablation studies according to each instrument. Table 12 and 13 show the full results of Table 4 (Section 4.1), and Table B.4 shows the full instrument-wise results of Table 4.2 (Section 4.3).
Table 12: Intervention with Multiple Directions - Full Results (1)
Individual Success Rate [%]
start with drums start with bass start with guitar start with piano
Method bass guitar piano avg. drums guitar piano avg. drums bass piano avg. drums bass guitar avg. total
unconditioned $29.0$ $4.3$ $0.2$ $11.1$ $17.7$ $10.6$ $1.4$ $9.9$ $18.8$ $18.4$ $1.9$ $13.0$ $11.9$ $10.4$ $9.1$ $10.4$ $11.1$
text $58.4$ $2.6$ $5.2$ $22.0$ $45.7$ $2.8$ $2.9$ $17.1$ $45.5$ $67.0$ $4.8$ $39.1$ $42.9$ $50.4$ $6.6$ $33.3$ $27.8$
SMITIN $48.3$ $16.1$ $01.6$ $22.0$ $25.4$ $23.6$ $04.1$ $17.7$ $41.1$ $49.0$ $3.3$ $31.1$ $30.2$ $35.7$ $26.2$ $30.7$ $25.3$
Table 13: Intervention with Multiple Directions - Full Results (2)
Simultaneous Success Rate [%] ($\uparrow$) FADL-CLAP ${}_{\text{L-CLAP mus}}$ ($\downarrow$) Similarity ($\uparrow$)
Method drums bass guitar piano avg. drums bass guitar piano avg. drums bass guitar piano avg.
unconditioned $0.0$ $0.0$ $0.2$ $0.2$ $0.1$ 0.442 $0.517$ $0.280$ $0.440$ $0.419$ $0.884$ $0.580$ $0.844$ $0.817$ $0.781$
text $0.3$ $0.0$ $0.5$ $0.9$ $0.42$ 0.364 $0.426$ $0.376$ $0.312$ $0.369$ $0.847$ $0.565$ $0.589$ $0.655$ $0.664$
SMITIN $0.1$ $0.4$ $0.5$ $3.9$ $1.225$ 0.467 $0.562$ $0.3$ $0.452$ $0.445$ $0.837$ $0.549$ $0.79$ $0.763$ $0.734$
Table 14: Effects on Number of Probing Data - Full Results (shows full-set success rate)
Number of Paired Data Success Rate* [%] ($\uparrow$) FADL-CLAP ${}_{\text{L-CLAP mus}}$ ($\downarrow$) Similarity ($\uparrow$) Probes Max. Acc [%] ($\uparrow$)
drums bass guitar piano avg. drums bass guitar piano avg. drums bass guitar piano avg. drums bass guitar piano avg.
$n$ = 10 $28.3$ $42.9$ $48.4$ $4.5$ $31.0$ 0.344 $0.267$ $0.445$ $0.373$ $0.357$ $0.855$ $0.928$ $0.920$ $0.935$ $0.909$ $88.6$ $85.3$ $73.3$ $69.0$ $79.1$
$n$ = 100 $25.0$ $48.9$ $45.0$ $8.3$ $31.8$ 0.368 $0.289$ $0.406$ $0.336$ $0.349$ $0.866$ $0.920$ $0.929$ $0.941$ $0.914$ $91.6$ $87.8$ $78.3$ $72.8$ $82.6$
$n$ = 500 $24.5$ $41.0$ $38.2$ $5.8$ $27.3$ 0.360 $0.280$ $0.395$ $0.330$ $0.341$ $0.865$ $0.923$ $0.934$ $0.940$ $0.915$ $93.6$ $88.8$ $81.0$ $74.1$ $84.4$
$n$ = 1000 $25.1$ $37.6$ $34.1$ $5.5$ $25.5$ 0.357 $0.273$ $0.400$ $0.331$ $0.340$ $0.861$ $0.927$ $0.930$ $0.940$ $0.914$ $94.3$ $89.1$ $81.4$ $74.7$ $84.8$
[…]
Appendix C Analytical Insights from Probing
C.1 Exploring a wider range of probing tasks
To expand our understanding of MusicGen’s capabilities beyond recognition on instrument and realism, we conducted probing on a set of more general music understanding tasks. These include music tagging (MTT) (Law et al., 2009), genre classification (GTZAN) (Tzanetakis & Cook, 2002), key detection (GS) (Knees et al., 2015), and emotion recognition (EMO) (Soleymani et al., 2013). We adopt appropriate regression models for probing for these tasks: multinomial logistic regression for multi-class classification tasks, multiple logistic regressions for multi-label classification tasks, and linear regression for regression tasks. For these tasks, we set the input duration as 29.0 seconds, and let the probes take the last time step activation output for classification.
Table 15 presents the probing results of generative music transformers, including various configurations of MusicGen and a comparative model, Jukebox (Dhariwal et al., 2020). We include objective results reported in (Castellon et al., 2021) for Jukebox probing. Jukebox, with its larger number of parameters and activation dimension size, generally shows superior performance. However, our primary interest lies not in the superiority of one model over another but in demonstrating MusicGen’s comprehensive understanding of various musical aspects, indicative of its versatility for ITI.
The insights gained from these probing results reveal MusicGen’s proficiency across a spectrum of musical elements, extending from instrument recognition to more complex attributes like genre and emotion. This opens up possibilities for controlled and targeted music generation, where MusicGen can be guided to produce outputs aligning with specific musical characteristics or themes.
Table 15: Probing generative music transformers.
Dataset MTT GTZAN GS EMO
Task Tagging Genre Key Emotion Num. Param.
[…]
Jukebox $91.5$ $41.4$ $79.7$ $66.7$ $72.1$ $61.7$ $69.9$ 4.8K 5B
[…]
C.2 Full Probing Results
We disclose heatmaps of probing classifier performances on every probing task performed in this paper. Heatmaps from different model configurations are normalized in the same scale for each task.